OpenStudy (ria23):

y^2+4y-8x+4=0 (x-8)+(y+2)^2=-4

OpenStudy (ria23):

There was only one x... I got kinda confused. ;c

OpenStudy (mathstudent55):

What are you asked to do?

OpenStudy (ria23):

Put it into standard form

OpenStudy (jtvatsim):

@Ria23 You will need a different standard form this time. The last one we did had x^2 and y^2... since this has only one x (like you noticed) this will be a parabola standard form.

OpenStudy (ria23):

>~< Why are there so many, oh my goodness...

OpenStudy (solomonzelman):

\(\large\color{black}{ y^2+4y-8x+4=0 }\) \(\large\color{black}{ y^2+4y+4=8x }\) \(\large\color{black}{ (y+2)^2=8x }\) then divide by 8 on both sides.

OpenStudy (solomonzelman):

\(\large\color{black}{ x=\frac{1}{8}(y+2)^2 }\)

OpenStudy (jtvatsim):

@Ria23 haha, I know. :)

OpenStudy (solomonzelman):

looks like a horizontal parabola.

OpenStudy (ria23):

Why did the 8x get moved? .-.

OpenStudy (solomonzelman):

I added 8x to both sides.

OpenStudy (ria23):

And then the last step, when yhu divide it, why is it 1/8 and not just an 8 in the front?

OpenStudy (solomonzelman):

I divided both sides by 8.

OpenStudy (jtvatsim):

@Ria23 I'm looking up some reference tables for you, are you in an Algebra II class?

OpenStudy (ria23):

Pre-calculus

OpenStudy (jtvatsim):

OK

OpenStudy (ria23):

@jtvatsim in the previous question yhu helped me with, was that a hyperbola?

OpenStudy (jtvatsim):

No that was a circle. But a hyperbola has a similar equation just one of the "squares" is negative and on is positive. I think I found something, are the questions you are working on from a lesson on conic sections?

OpenStudy (ria23):

Yes indeed. c:

OpenStudy (jtvatsim):

OK, I'll upload it in a second

OpenStudy (ria23):

Ok. c:

OpenStudy (jtvatsim):

I have to make some modifications to it so that it makes sense.

OpenStudy (ria23):

Alright. c: Thank yhu so much c:

OpenStudy (jtvatsim):

no worries, almost done :)

OpenStudy (jtvatsim):

there were a bit more of these things than I thought... :)

OpenStudy (ria23):

xD Yhu don't have to do this. But it's really nice of yhu to do it, thank yhu, if yhu don't want to, yhu don't have to. Haha

OpenStudy (jtvatsim):

Got it... all done! :)

OpenStudy (jtvatsim):

and the original version, more detailed.

OpenStudy (ria23):

Thank yhu. n.n